OpenStudy (anonymous):

Please help!! 1/4 x 1x3 + 1/2 x 1/6

OpenStudy (mathmale):

Please help me and other potential helpers by including the instructions for this problem. Need clarification: when you use 'x' in this statement, is 'x' a variable, or do you mean 'multiplication?'

OpenStudy (mathmale):

How you present a problem is extremely important. I think you mean\[\frac{ 1 }{ 4 }x^3; \]I can't figure out the second term. Present your problem using Draw (below), if at all possible.

OpenStudy (mathmale):



OpenStudy (anonymous):

here is a new problem \[\frac{ 3 }{ 4 } \times \frac{ 1 }{ 5 } + \frac{ 1 }{ 2} \times \frac{ 3 }{ 10 }\]

OpenStudy (mathmale):

If you must indicate multiplication, please do not use the character 'x'. Instead, us the character '*' OR use parentheses to indicate multiplication. I think that by writing 1/2 x 1/ 6 you actually meant (1/2)(1/6), or \[\frac{ 1 }{ 2 }*\frac{ 1 }{ 6 }\]

OpenStudy (mathmale):

Please let's finish your original problem before moving on to the next. I'd like for you to respond to my suggestions. Please take your original expression and re-write it for clarity.

OpenStudy (mathmale):

Sorry if you feel I'm asking too much of you. However, we all need to know how to present math questions clearly enough so that it's not necessarily for anyone to ask you what you meant by this or that.